OpenStudy (anonymous):

Under water, pressure increases 4.3 pounds per square inch (psi) for every 10 feet you descend. This is represented by the equation p = 0.43d + 14.7, where p is the psi and d is the depth in ft. Find the pressure at 200 feet. A. 215.13 psi B. 860 psi C. 100.7 psi

OpenStudy (anonymous):

p = 0.43d + 14.7 p = 0.43 X (200) + 14.7 p = 86 + 14.7 p = 100.7!!!!!